The length of a rectangle is a = 3 cm less than twice the width. Express as an integer the maximum width of the rectangle when the perimeter is less than 80 cm.


This is what I have:
80=6w-6
+6 +6
86=6w
------
6 6
14.3333333=w
14.33=w

Answer by DrBeeee(684)

Always check your answer. If the width is 14.33 then the lenght is 28.66 - 3 or 25.66, then the perimeter is 2*(14.33+25.66) or 80. So your answer is correct.
Let w = the width and
Let a = length
Then, from the problem statement we have
(1) a = 2*w - 3
The perimeter, p, is the sum of the 4 side or
(2) p = 2*a + 2*w or
(3) p = 2*(2*w - 3) + 2*w or
(4) p = 4*w - 6 + 2*w or
(5) p = 6*w - 6
Using the maximum value of p as 80 we have
(6) 6*w - 6 <= 80 or
(7) w <= 86/6 or
(8) w <= 14.33...
Answer: The largest integer value of the width is 14 cm.